

********************************************************************
UARS WEEKLY STATUS REPORT - 24 AUG 2001
********************************************************************
>>> Latest on UARS Situation (from the UARS Project Science Office) <<<
August 24, 2001
UARS Colleagues:
This message is being sent to bring you up to date on the latest
news regarding the UARS Mission:
UARS INSTRUMENT DATA COLLECTION - Instrument science observations
are scheduled to cease on September 24th. The instruments will
be placed in survival modes pending further decisions.
UARS CENTRAL DATA HANDLING FACILITY (CDHF) - The UARS CDHF is in
the midst of final processing and data archive. It is hoped that
NASA Headquarters will support this effort through November 2001,
by which time most of the archival will have taken place. The
archive will primarily be taking place on Digital Linear Tapes
(DLTs), however, most of the data (L0/L2/L3) will also have been
moved to the new rehost computer system by then. Practically
all the L3 (L2 for HALOE) data will have been moved to the DAAC
by this time as well.
UARS PROPULSION SYSTEM - The backup thrusters have been burned
successfully. There may be some more test or calibration burns
in the next two weeks, however, NASA Headquarters upper
management has informed the project that they should not take any
action that is "irreversible." These thrusters are crucial to
controlling the altitude of the satellite and will be needed when
the satellite is retrieved or decommissioned.
UARS MISSION OPTIONS - NASA Headquarters upper management is
considering several possibilities regarding the UARS, including
the possible shuttle retrieval. The "Safety and Mission
Assurance" part of NASA is to give their recommendations in early
September on the decommissioning of the UARS satellite. It is
still not clear which option will be implemented. Presently,
there are ZERO dollars in the FY02 budget for UARS instrument
observations.
Thanks for bearing with the UARS Project through this difficult
situation. Your words of support are greatly appreciated. We
will continue to inform you of any further developments/changes
in the UARS mission status.
Sincerely,
Charles Jackman and Anne Douglass
UARS Project Scientist UARS Deputy Project
Scientist
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
Since resuming operations on Aug. 12, after the unexpected TMON power
off, SUSIM has completed its daily solar scans this past week during
periods of predicted UARS telemetry.
We are preparing for September operations.
The SUSIM team would like to thank all of those who have supported
UARS making it possible to continue for another month.
***********
ACRIM
***********
The TMON event of 8/17/01 has re-activated the ACRIM2 experiment. Good
total solar irradiance data has been produced by the instrument for days
8/18-20/01 thus far. A multi-sensor comparison will be conducted in the
next few days to determine whether the 3 1/2 month down time affected
sensor calibrations. Glad to be back in business for the last month or so
of the UARS Mission.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha). Currently, PEM has requested GSFC assistance with compiling for
the DEC Alpha since the GSFC executables run on the Alpha and SwRI's compiled
versions do not.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
On 23 August 2001 the MPG participated in the weekly UARS
Technical Team Meeting. This series of meetings is being
held to discuss contingencies to be implemented in the
event that FY 2002 funds are insufficient to continue
collection of science data by the UARS instruments.
The processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 11 August 2001 (Orbit 54223, GMT Day 223) through 17 August
2001 (Orbit 54316, GMT Day 229).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 224/1233 54233 Power On
HALOE 225/2353 54256 Power On
HALOE 228/2356 54301 Power Off
WINDII 228/2356 54301 Power On
MLS 229/0141 54302 Power On
ACRIM 229/0513 54304 Power On
SOLSTICE 229/0513 54304 Power On
WINDII 229/0550 54304 Power On
WINDII 229/1339 54309 Power Off
The beta angle ranged between the angles of 9.7 degrees to 31.3 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 46 degrees on 25 August 2001 (DOY 237).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.9 to 2.2 degrees C. Additional battery performance data for each
battery on 10 August 2001 and 17 August 2001 are provided in Appendix D.
The clock error ranged from Ð9.3 to 2.9 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 03
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 08
¥ ACRIM = 02
¥ SUSIM = 02
¥ PEM = 02
¥ HALOE = 03
¥ WINDII = 05
¥ SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/228 show the observatory data loss to be 303 hours, 06
minutes, 40.256 seconds. This shows a loss of 06 minutes and 31.008 seconds
since the last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
The increase of 06 minutes and 31.008 seconds occurred over three (3) days:
· 5 minutes and 49.392 seconds on Day 222 due to a Late AOS TTR# 14170
· 2.768 seconds on Day 223 due to a Data Hit
· 38.848 seconds on Day 224 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS FOT experienced a low power situation on DOY 229 at approximately XXXX
GMT. This was caused due to both WINDII and MLS being turned on with not
enough power margin to support them. Subsequently, both TMONs 1 & 2 were
tripped which turned off all instruments that were scheduled to be on (MLS,
WINDII, SOLSTICE, SUSIM, ACRIM, HRDI and PEM). The FOT immediately caught the
TMON trip and began to turn all instruments back on as soon as possible.
One good note to the tripping of the TMONs was that when ACRIM was turned back
on, it came back to life and showed good telemetry. ACRIM PIs were notified
and they have begun to take a look at the data to verify if ACRIM is once again
taking good science. Unfortunately, they may not get to take data for very
long, since all science data processing will be turned off on 27 August 2001.
MEETINGS
The UARS FOT attended a UARS Technical Meeting on Thursday, August 16, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 11 September 2001 (DOY 254). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 August 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
54211 223/0033 01-504 OBC Expanded Ephemeris Tolerance
54212 223/0132 01-505 OBC Normalized Ephemeris Tolerance
54219 223/1414 01-506 SUSIM Power ON
54224 223/2111 01-434 ACRIM Power ON
54233 224/1148 01-506 SUSIM Power ON (complete)
54233 224/1233 01-507 PEM Power ON (ZEPS & HEPS)
54256 225/2353 01-509 HALOE Power ON (start)
54260 226/0727 01-510 PEM Day-Only Operations
54262 226/1011 01-511 HALOE Power ON (complete)
54274 227/0548 01-512 NBTR NOOPS D-227
54289 228/0549 01-513 NBTR NOOPS D-228
54301 228/2356 01-515 HALOE Power OFF
54301 228/2356 01-516 WINDII Power ON (start)
54301 229/0034 01-516 WINDII Power ON (complete)
54302 229/0141 01-517 MLS Power ON (start)
54303 229/0329 01-514 NBTR NBTR PLAYBACK
54304 229/0513 01-514 NBTR RTS-58 ACTIVATE
54304 229/0513 01-518 ACRIM Power ON
54304 229/0513 01-519 SOLSTICE Power ON
54304 229/0550 01-514 NBTR NOOPS D-229
54304 229/0550 01-516 WINDII Power ON (start)
54305 229/0654 01-516 WINDII Power ON (complete)
54309 229/1339 01-520 WINDII Power OFF
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 August 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
OBSVLBEON Y-LO Load Bus Volts at EON 54259
OBSBATVEON3 Y-LO Battery 3 EON Voltage 54259
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 54259-60
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
__
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 August 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 August 2001
10 August 2001 - GMT Day 222 Beta = 5.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 12.0 0.000 0.0
2 +16.8/-67.2 +5.15/+3.87 +28.8/-12.8 25.4 1.040 20.5
3 +22.4/-56.0 +4.19/+2.92 +28.0/-12.8 25.4 1.030 20.5
17 August 2001 - GMT Day 229 Beta = 31.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-95.2 +5.47/+3.23 +28.0/-16.8 24.5 1.020 19.3
3 +22.4/-50.4 +4.51/+2.60 +27.2/-16.8 24.5 1.020 18.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 August 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-046 FOT 54211 223/0033 14170 Generic Late Acquisition
01-046 FOT 54278 227/1235 14170 Generic Late Acquisition
01-050 FOT 54303 229/0329 N/A TMONS 1&2 tripped - Low Power
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-050 FOT 54303 229/0329 N/A TMONS 1&2 tripped - Low Power
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-045 FOT see list August 01 14170 Generic Late Acquisition
SEVNTFW01/33